A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 February 2018 Version 1.02

Title word cross-reference

A [Ise78]. A_{0} [RT79]. C^{2} [Wat79]. Δ^{2} [Col78]. g [CG77]. M [BE79]. n [Dah77, vDdR76]. q [SD79]. tan(ζ − kπ) + tanh ζ = 0 [SP79]. × [Bul76, Goo78b, Goo78a, Wat78]. x coth x = αx^{2} + 1 [SB76b].

-acceptable [Ise78]. -dimensional [Dah77, vDdR76]. -polynomials [SD79]. -splines [CG77]. -stable [RT79].

0355-2713 [Goo77a].

1 [Lat75, Neu77, SB76a, WA75]. 18 [Goo78a]. 1975 [dG77]. 1977 [Goo78b].

1978 [Kau79].

2 [Ano76c, BS76, WA76]. 22 [Wat78]. 25 [Ano75a].

3 [SB77].

80 [Goo78b].

951 [Ano76a, Goo77a]. 951-666-069-X [Ano76a]. 951-666-086-X [Goo77a]. 99 [Goo77b].

A-stable [Alt78]. A.D.I. [AM77]. accelerating [Lev75]. acceleration [Gen77]. acceptable [Ise78]. accuracy [AM77]. Acta [Ano75a, Ano75c, Ano76a, Goo77a]. Adams [RT79]. Adams-type [RT79]. Adaptation [Ano75a]. adaptive [vDdR76]. adjoint [AH78, PL77]. Agricultural [Bul76]. Aitken [Col78]. algebra [WA75, WA76]. algebraic [BS78c]. algorithm [Abb78, BE79, BS76, De 76, Die75, Die77, MC75, Mik77, MHH75, Nou77, Pin76, SB76a, SB77, dB78, vDdR76]. Algorithms [O’S78, CG77]. allocation [Lat75]. alternating [AGH79]. amplitudes [Ano76c, CGP76]. Amsterdam [Wat78]. analysis [BM76, FK75, GMJ76, Rob77, Goo78a]. Analytic [Dev76]. Analytical [AH78, Let76a, SB76b]. annoted [Neu77]. Appl [Ano76c]. Application [De 79, Lon75, Dey78]. Applications [SB77, Ans76, DDG76, Dey77, Goo78b, Pie75, PD76, WA75, WA76]. applied [BF78, Mee75, NS75]. approach [Abd79, Hen78, Pes75, Pic76, SB78]. approximants [Bre76, CW79, GMJ76, Rob77, Sta79]. approximation [Col78, Dun79a, Dun79b, KM78, Lit78, Lon75, Neu78, Pau76, Ray77, Sid75b, Sid76]. approximations [Dev76, Ise78, Luk76, Oli77, Sas76]. arbitrarily [Ano75a]. areas [Kau79]. arising [AH78]. aspects [AS78]. assignment [Raz78]. associated [SD78]. Asymptotic [BS78b, Ans76]. asymptotical [SD77]. automatic [DP76]. averaging [Mel78]. AY=F [HMW77].

B.V [Bul76]. Baker [Pin76]. balances [De 79]. bandwidth [BE79]. based [BM76, CG77]. Basic [SB76a]. Beckmann [Kau79]. behaviour [Ans76]. Berlin [Goo78b, Goo78a, Kau79]. Berry [GV78]. Best [Lit78, Sid76, Dun79b, Lon75]. between [CK75, Rah78]. Bibliography [Ein79, BS78a, DP76, De 75a, DDG76, KM78, Kau79, Neu77, Pie75, PD76, Teu76]. bifurcation [Abb78]. Birkhoff [Let76a]. blz [Wat78]. Bonn [Kau79]. Book [Ano75a, Ano76a, Bul76, Goo77a, Goo77b, Goo78b, Goo78a, Goo78c, Kau79, Wat78, dG77]. Borel [BS76]. Born [Col78, Ray77]. boundary [Aga79a, Gra77, JS78, Wat78, WC79]. boundary-value [JS78]. bounded [Mav79, Pau76]. bounds [AH78]. branch [Ano76c, CGP76]. Bromwick [Lev75]. buckling [BF78].

calculating [Mik77, Pin76, dB78]. Calculation [DP77, HM79]. Centre [Bul76, Goo77b, Wat78]. Centrum [Goo77b, Wat78]. certain [Abb78, BS78c, MV79]. change [RM78]. characteristics [CG77, SD76]. Chebyshev [Dun79a, Luk76, Oli77, Sal77a, Sid75a]. choice [Dan76]. Choosing [CL77]. class [Ver77]. classified [Kau79]. Clenshaw [BP75]. Clique [MC75]. clock [FK75]. closing [De 79]. cm [Bul76, Goo78b, Goo78a, Wat78]. coefficient [De 75c]. coefficients [AM77, DP77, JC78]. collocation [AR79]. combinations [Ise78]. combined [Dev76]. Company [Goo78c]. comparative [Hen78]. Comparing [vdHSV79]. comparison [CK75, NH76, Rob79]. Compiled [Kau79]. complex [SW77]. complexity [Hou78, Mee75]. component [BS76]. compound [GV78]. compressed [BF78]. Comput [Ano76c]. Computation [Bre76, Dah77, Ebe77, Eva77, LR79, Let76b, PB75, Sid75a, Abe78, CW79, Mik75, SW77]. computational [AS78, Mee75, Neu77, PP75]. Computations [GJ79]. Computer [Ano76a, Goo77a, Gre77, GC78]. Computing [Ano75a, Ano75c, Han79, Ano76c, Bul79, CGP76, Cla75]. condition [Raz76]. conditions [SD76]. connected [GJ79]. consistent [Pes75]. constants [GJ79]. constraints [De 79, KL75]. Construction [Lew79, GM77, Let78, PL77]. Contents [Ano75b, Ano76b, Ano77b, Ano78]. continuation [Dev76]. continued [BS78b, Bre76, Mik77, MO78, SB78, dB78]. Continuous [Sny79]. control [Alt78, Ano76a, FK75]. controller [Goo77a]. convergence [Gen77, Lev75, Sac79]. convergent [HMW77]. coordinate [Mel78]. correct [RZ76]. correctors [RT79]. Coulomb [Pic76]. coupled [Dek77, Dey78, O’S78]. covariant [Rob77]. cover [Kau79]. Crane [Goo78a]. credibility [DDG76]. Cubature [PH75]. cubatures [Wil78]. cube [vDdR76]. cubic [Die77, DP77, Sas76]. cumulatively [Mav79]. Current [Ano77a]. Curtis [BP75]. curves [Joh77]. customers [Lat75]. CUSUM [RM78]. cuts [Ano76c, CGP76]. cyclic [De 75c].

D [Goo78c]. data [Die75, Ein77]. decision [Pes75]. definite [De 76]. definiteness [Spi77]. Degree [Wil78, PH75]. delays [Raz78]. density [SD76]. derivation [De 76]. derivative [GM77, vdHSV79]. derived [Pin76]. design [Goo77a]. detecting [RM78]. detection [MC75]. determinant [Mik75]. determination [Abb78, Nou75, Nou77, SS76]. DFP [RZH75]. DFP-method [RZH75]. difference [JS78, Sal77b, Sas76]. different [Bul79, Cla75, IKSN79]. differential [Aga79b, Alt78, Ano75c, AH78, Dek77, EF77, GM79, Hou78, Lit78, MV79, O’S78, SW77, Wam76, Goo78b]. differentiation [Die75]. differing [Gre77]. diffusion [EG79]. dimensional [AGH79, Dah77, GM78, Sas76, vDdR76]. dimensions [GM77, vdHSV79]. directed [MC75]. direction [AGH79]. Dirk [Kau79]. discontinuous [Sny79, WC79]. discrete [DP77]. discretizing [AH78]. disk [Rah78]. divergent [BS76, SB76a, SB77]. divisibility [Goo77b]. DM [Goo78b, Goo78a, Kau79]. Documentation [Bul76]. domains [AR79]. double [Let76b, Rob77]. ducts [CB76]. dynamics [Dey77, Dey78].

E.A.D.I. [AGH79]. Economics

[Kau79, dG77]. Edited [Kau79, Wat78]. Editorial [PW75, Wuy79]. Editors [Kau79, Goo78b, PW77, dG77]. efficient [Abb78, Spi77]. eigen [SD76]. eigen-value [SD76]. eigenfunctions [PL77]. eigenvalue [AH78]. eigenvalues [BO77, Eva77, HM79, SW77]. eigenvectors [Eva77]. elastica [BF78]. elements [Bul79, Cla75, Sac79]. eleven [PH75]. elliptic [AH78, AGH79, GM77, GM79, Hou78]. engineering [Raz76, WA75, WA76, Goo78c]. Enschede [Goo78b]. equality [De 79]. equation [BW78, HMW77, NH76]. equations [Alt78, AM77, Ans76, AH78, AS78, BS78c, CK75, Dek77, Dey77, EF77, GM77, GM79, HM79, Hou78, Lit78, MV79, O’S78, Sas76, Ver77, Wam76, vdHSV79]. equiconvergence [GJ79]. Erkki [Ano75a]. Erratum [Ano76c]. error [Oli79, Sal77b]. errors [Oli77]. Esseen [GV78]. estimating [Abd79]. estimation [Ano75c]. evaluation [BDG75, Ein79, Oli79]. evolution [GC78]. exact [Abe78, SB76b]. expansions [Luk76]. experiment [KL75]. experimental [Die75]. experiments [RZH75]. Explicit [Sta79]. exponential [Dun79a, KM78, Sid75b]. exponential-polynomial [Dun79a]. exponentially [Ise78]. extension [BP75, HM79]. extrapolated [AGH79].

F [Goo77b, Wat78]. feature [BO77].

Feynman [Ano76c, CGP76]. Fifth [CG77]. fig [Goo78b]. Finite [Sas76, Ano75a]. fitted [Ise78]. fitting [Die77]. five [BE79]. five-term [BE79]. Fixed [Wat79]. flow [CB76, Rah78]. fluid [Dey77, Dey78]. form [Sal76, WA75, WA76]. formula [JC78, Nou75]. formulas [DS77, Let77, PH75, Sal77b, Sta79]. Fortran [BW78]. Fourier [DP77]. fourth [AM77]. fraction [MO78]. fractions [BS78b, Bre76, Mik77, SB78, dB78]. function [BS78b, Lon75, RZH75, SD78, Sid76, Wam77]. functional [Ans76]. functionals [Pau76]. functions [BDG75, Die75, Die77, DP77, Kle79, Let76a, Let77, MHH75, RM78, Sid75b]. fuzzy [De 75a].

G [Goo78b, dG77]. game [Goo77a]. games [Goo78b]. gamma [BS78b]. Gauss [Let78]. Gaussian [LR79, Sal76, Wil78]. Gaussian-type [Sal76]. general [AR79, BE79, SD79]. generalization [MO78]. Generalized [Dek77, Miz75a, Miz75b, SB78, Chi79, dB78]. generating [Ein77]. graphs [MC75]. Gustafsson [Ano76a].

H [Bul76, Goo78b, Kau79]. Haar [Van79]. Hagedorn [Goo78b]. Hämäläinen [Goo77a]. Hamiltonian [SD76]. Hausman [Kau79]. Heidelberg [Kau79]. Helsinki [Ano76a]. Hemker [Wat78]. Hermite [Miz75a, Miz75b, Neu78]. Hessian [Spi77]. High [AM77, Axe75]. High-order [Axe75]. Higher [JS78]. Hilbert [Pau76]. Hill [HM79, PL77]. Hopscotch [EG79, GM79, GM77, GM78]. Householder [Dan76]. hyperbolic [Dek77]. hyperosculatory [Sal77a].

IBMSSP [CK75]. ill [Kle79]. ill-posed [Kle79]. illustrations [SB76a]. Implementation [EG79]. Implicit [Mel79, AGH79]. improved [CG77]. improvement [Nou77]. IMSL [CK75]. index [Ano77a]. inequalities [Ein77, Han79]. infinite [Goo77b]. information [Pes75]. inhomogeneous [CD78]. inner [Pau76]. input [Raz78]. Institut [Kau79]. Institute [dG77]. Integer [Kau79]. integrable [Sid75b]. Integral [PP75, Lev75]. integrals [BDG75, Let76b, PB75]. integration [DP76, Die75, EF77, Rob79, Ver77, Wuy75, vDdR76]. integrators [vdHSV79]. interactions [Gre77]. interdependent [BM76]. interior [Han79]. International [dG77]. interpolating [Neu78]. Interpolation [Joh77, BM78, De 75b, MHH75, Sal77a]. Interval [BS78a, BM78]. Interval-mathematics [BS78a]. intervals [Van79]. introduction [Goo78a, KL75]. invariant [Pau76]. Inverse [Sal77a]. inverses [DS77]. inversion [Hen78, JC78, Lev75, Lon75, Pie75, PD76, Sid76]. ISBN [Ano76a, Goo77a]. ISSN [Goo77a]. Italy [dG77]. iteration [IKSN79, Nou75]. iterative [Dey77, Dey78].

J [Ano76c, Goo78b, Goo78a]. Jacobi [IKSN79, Raz76, SD77, SD78]. junction [CD76]. June [dG77]. Jussi [Ano75c].

Karel [Goo78c]. Karup [De 75b]. kernel [Dev76]. Knobloch [Goo78b]. known [De 77]. Kohonen [Ano75a]. Künzi [Kau79]. Kutta [Dek77, Ver77].

laminar [CB76]. Laplace

[Dah77, Hen78, JC78, Lev75, Lon75, Pic76, Pie75, PD76, Sal77a, Sid76]. Laplace-transform [Pic76]. Least [De 79, JC78, De 75c, DS77, Die77]. least-squares [Die77]. Lecture [Kau79, dG77]. Legendre [Let78]. Lemoine [Goo78a]. Letter [PW77]. level [AGH79]. Levinson [De 76]. linear [Ano75a, Ano75c, Ano76a, BE79, BO77, CK75, De 76, De 79, Ein77, EF77, FK75, Han79, MV79, NH76, NS75, Pau76, SW77, WC79, WA75].

LINSYS [CK75]. local [Joh77]. Locally [GM78]. Longman [Pin76]. look [Bul79, Cla75]. loop [RM78]. Low [RT79]. Low-order [RT79].

M [Goo78a, Kau79, Lat75, Lat75]. M/M/1 [Lat75]. Machinery [Ano75a, Ano75c]. maintenance [Spi77]. maps [Wat79]. Marchesini [dG77]. Markov [Teu76]. Massachusetts [dG77]. masses [Gre77]. Math [Ano76c]. Mathematical [BM76, Goo77b, Kau79, Wat78, CD76, CD78, dG77, dG77]. Mathematics [Ano75a, Ano75c, Ano76a, Goo77a, BS78a, Goo78c]. Mathematisch [Goo77b, Wat78]. matrices [AH78, Dan76, SD77, SD78, WA76]. matrix [De 75c, De 76, Eva77, HMW77, IKSN79, Mik75, SD76, Sta79, WA75, WA76]. matrix-valued [Sta79]. Matti [Ano75a]. mechanics [PP75]. media [IKSN79]. mesh [Die77]. method [Abe78, Ano76c, AR79, BW78, BF78, CGP76, Col78, De 75b, De 75c, Gen77, Goo78a, HM79, HMW77, Mel78, Nou77, Sac79, SW77, NS75, RZH75, Goo77a]. methods [Alt78, AM77, AGH79, Axe75, Bul79, Cla75, CL77, Dek77, Dev76, Goo78c, GM77, GM78, GM79, Hou78, JS78, Lit78, Spi77, Ver77, Wam76, WC79]. Minimal [Sal77b]. minimization [RZH75]. Mitter [dG77]. mixed [GM77, vdHSV79]. mixing [Goo77b]. model [BS76, CD76, CD78]. models [Abd79]. Modified [Let77, LR79, Lew79]. moments [LR79, Lew79]. Muller [BW78]. multidimensional [Gen77]. multipoint [Aga79a]. multistep [EF77]. multivariable [Ano76a]. Multivariate [BM78].

near [Sal77a]. near-Chebyshev [Sal77a]. nearby [Let77]. Newton [Spi77]. Newton-type [Spi77]. No [Ano75a, Ano75c, Ano76a, Goo77a]. nodes [Sal77a]. non [Abd79, Ano75c, CW79, De 79, FK75, NH76]. non-linear [Ano75c, De 79, FK75, NH76]. non-normal [CW79]. non-stationary [Abd79]. Nonlinear [CD76, CD78, Mav79, Wam76, WA75, WA76, Aga79b, BW78, Dey77, Dey78, Gen77, Goo77a, Wuy75, Bul76]. nonsymmetric [NS75]. normal [CW79]. normalised [BE79]. note [BDG75, Wuy79]. Notes [Kau79, dG77]. Nourein [Nou77]. numbers [Abe78, GJ79]. Numerical [CB76, Lev75, RZH75, WC79, Wuy75, Aga79a, AGH79, AS78, BDG75, Ein79, EF77, EG79, Gra77, Hou78, Lit78, Pie75, PD76, Rah78, Rob79, RZ76, Sac79, Sny79, SU75, Ver77, Wat78, vDdR76]. numerically [Spi77].

occurring [Ano75a]. ODE’S [CL77]. Oja [Ano75a]. Ökonometrie [Kau79]. Olsder [Goo78b]. one [Alt78, GM78, Sas76]. one-dimensional [GM78, Sas76]. one-step [Alt78]. open [RM78]. Operations [Kau79]. operators [PL77]. Optimal [Lat75, Ano76a, Ebe77, FK75, Pau76, Goo77a]. optimization [Han79, Bul76]. optimizing [Raz76]. Orava [Ano75c]. order [Aga79b, AM77, AH78, AGH79, Axe75, CG77, JS78, O’S78, RT79, Ano75a]. ordinary [Alt78, EF77, MV79, O’S78, Wam76]. orthogonal [Chi79, SD79, SS76]. orthogonality [Sal76]. oscillating [BDG75, PB75]. oscillatory [EF77]. osculatory [Sal77a].

P [Goo77a, Goo78b, Kau79, Wat78, Goo78a]. packages [CK75]. Padé [Bul79, Bre76, Cla75, CW79, Dev76, Ise78, Pin76, Sid75a, Sta79]. Padua [dG77]. pages [Kau79, Bul76, dG77]. parabolic [AM77, AR79, Axe75, AS78, GM77, Sas76, Ver77, vdHSV79]. parameter [Bul76]. parameters [Abd79, KL75]. Part [BS76, Neu77, SB76a, SB77, WA75, WA76]. partial [AH78, GM79, Hou78]. particles [Gre77]. patterns [Ano75a]. PECE [CG77]. pentadiagonal [Mik75]. performance [CK75]. periodic [Aga79b, MV79, PL77]. perturbation [BF78, FK75]. perturbations [Mel79]. Perturbed [Dey77, Dey78]. Ph. [Bul76]. PI [Ano76a]. PI-control [Ano76a]. pipes [CB76]. planar [PH75]. plane [Wil78]. planetary [GC78]. planetary-type [GC78]. plans [Ebe77]. Plotting [Wam77]. point [Gra77, Han79, JS78, Wat78, WC79]. points [Abb78, Wat79]. Poisson [GV78]. pole [Raz78]. poles [Let77]. Polynomial [WA76, Ano75c, Dun79a, Nou75, Nou77, Oli79]. polynomials [Chi79, Miz75a, Miz75b, Sal77a, SD79, Van79]. Polytechnica [Ano75a, Ano75c, Ano76a, Goo77a]. population [SU75]. posed [Kle79]. positive [De 76, Spi77]. Post [KL75]. Post-experiment [KL75]. power [Rob77]. powerful [RZH75]. pp [Ano76a, Goo78c]. prescribed [SD76]. Preservation [Goo77b]. price [Goo78b, Goo78a, Wat78]. priority [Lat75]. probabilistic [Pes75]. probability [Neu77]. problem [Abd79, Lap78, Pic76, Raz76, Sny79]. problems [Aga79a, AR79, AGH79, Axe75, EG79, FK75, Gra77, JS78, Kle79, WC79]. procedures [Pes75]. Proceedings [Goo78b, dG77]. processes [GV78, Teu76]. processing [Pes75]. product [Pau76]. program [BW78]. programming [Kau79, Mav79]. programs [Dun79b, Rob79]. proof [De 77]. propagation [Oli79]. Proposal [SU75]. pseudo [DS77]. pseudo-inverses [DS77]. published [Dun79b]. Publishing [Bul76, Goo78c].

quadratfrei [GJ79]. quadrature [BP75, JC78, LR79, Let76a, Let77, Let78, Sal76, SU75]. qualification [Pes75]. quantal [Pic76]. quantum [PP75]. quasilinear [AR79]. queue [Lat75]. queues [BM76]. quindiagonal [Eva77]. quintic [MHH75].

radiation [Ebe77]. radius [IKSN79].

Raimo [Goo77a]. rapidly [HMW77]. ratio [BS78b]. rational [Abe78, GMJ76, Lon75, Luk76, Sid76]. rectangle [Die77]. rectangular [De 75c, IKSN79]. rectilinear [Die77]. recurrence [Lew79]. recursive [Mik75]. refined [Col78, Ray77]. regerative [Goo78a]. region [IKSN79]. regions [PH75]. regularization [Kle79]. Reidel [Goo78c]. Rektorys [Goo78c]. related [Chi79, Goo77b, Kau79]. relation [Lew79]. relaxed [NS75]. Remark [Bul79]. reproducing [Dev76]. Research [Kau79]. result [De 77, MV79]. Review [Ano75a, Ano76a, Bul76, Goo77a, Goo77b, Goo78b, Goo78a, Goo78c, Kau79, Wat78, dG77]. Riccati [SW77]. risk [De 77]. rivers [CD76, CD78]. role [BO77]. rotating [Rah78]. rotation [Pau76]. rotationally [Rob77]. Rounding [Oli79, Oli77]. rules [LR79, Let78]. Runge [Dek77, Ver77]. Ruohonen [Ano75a].

S [dG77]. salesman [Lap78]. Scandinavica [Ano75a, Ano75c, Ano76a, Goo77a]. scheme [EF77]. schemes [Oli79, Pin76]. Schrödinger [HM79]. Science [Ano76a, Goo77a, Goo78c]. second [Aga79b, AH78, AGH79, JS78, O’S78]. second-order [O’S78]. selection [BO77]. self [AH78, PL77]. self-adjoint [AH78, PL77]. semi [BE79, Teu76]. semi-bandwidth [BE79]. semi-Markov [Teu76]. sensitivity [Oli77]. sequence [Sac79]. sequences [Gen77]. seriation [Lap78]. Series [Ano76a, Goo77a, Abd79, Ano75a, Ano75c, BS76, BS78b, CL77, Oli77, Rob77, SB76a, SB77, Sid75b]. set [Ano75a]. sets [De 75a]. shock [RZ76]. shuttle [BM76]. signs [Dan76]. simplified [Pin76]. simulation [Goo78a]. simultaneous [Nou75, Nou77, Sac79]. single [Raz78]. single-input [Raz78]. sixth [CG77]. size [Alt78]. smoothing [Die75]. Soft [Kau79]. software [Ein79]. solution [Aga79a, AGH79, AS78, BE79, BS78c, De 75c, Dey77, Dey78, EG79, Gra77, HMW77, Lit78, O’S78, Rah78, SB76b, SP79]. solutions [Aga79b, DS77, EF77, Sny79]. solvers [NH76]. Solving [Ein77, BW78, CK75, CL77, De 76, EF77, Wam76]. Some [AS78, Neu78, SD76]. SOR [NS75]. SOR-method [NS75]. space [GM77, Joh77, vdHSV79]. spaces [Pau76]. spectral [Abd79, IKSN79]. spectrum [SD77, SD78]. speed [RZ76]. spline [Die75, Die77, Kle79, Sas76]. splines [CG77, DP77, MHH75, Neu78]. Sprague [De 75b]. Springer [Goo78b, Goo78a, Kau79, dG77]. Springer-Verlag [Goo78b, Goo78a, Kau79]. square [Sid75b]. square-integrable [Sid75b]. squares [De 75c, DS77, De 79, Die77, JC78]. stability [CG77]. stabilized [Ver77]. stable [Alt78, RT79, Spi77]. State [Ano75c]. stationary [Abd79, Rah78]. statistical [Kle79]. statistics [Ano77a]. Sten [Ano76a]. step [Alt78, Ver77]. step-size [Alt78]. stepsize [CL77]. Steutel [Goo77b]. Stichting [Goo77b, Wat78]. Stieltjes [BS78b, Chi79, MO78]. Stieltjes-type [MO78]. stiff [Alt78, Wat78]. stochastic [RM78]. Stol [Bul76]. strongly [BDG75]. studies [CB76, Gre77, GC78, SU75]. study [CK75, Wat78]. subroutine [CK75]. summability [BS76]. summing [BS76, SB76a, SB77]. sums [Dun79a, KM78]. supplement [PD76]. surfaces [Die77]. survey [Dun79b, Mee75]. symmetric [BE79, De 76, Eva77, PH75]. Symposium [dG77]. system [Ano75a, BM76, Ein77, SD79]. Systems [Kau79, Aga79b, Alt78, Ano75c, BE79, CK75, De 75c, De 76, Dek77, Dey78, EF77, NS75, O’S78, RM78, Raz78, SW77, dG77, Ano76a].

tab [Goo78b]. table

[Bul79, Cla75, Pin76, Sid75a]. Tauberian [GJ79]. Taylor [CL77]. technique [EG79, RZH75, Sas76, Spi77]. techniques [Mee75, Wuy75]. Technology [dG77]. term [BE79]. terms [SD78]. test [Ein77, RM78]. Teuvo [Ano75a]. Th [Bul76]. theorem [Ans76, GV78]. theoretic [Goo77a]. Theory [WA75, Ano76a, BO77, DDG76, De 77, SB76a, dG77]. Three [AGH79, Ise78, Ver77]. Three-level [AGH79]. three-step [Ver77]. tidal [CD76, CD78]. time [Abd79, FK75, Raz78, vdHSV79]. time-delays [Raz78]. Toeplitz [De 76]. tools [PP75]. topics [Goo77b]. Tracts [Goo77b, Wat78]. transfer [RM78]. transform [Hen78, JC78, Lev75, Lon75, Pic76, Pie75, PD76, Sid76]. transformation [SW77, SS76]. transforms [Dah77, PP75, Sal77a]. translation [Pau76]. travelling [Lap78]. treatment [Ebe77]. triangle [Let76b]. trigonometric [Sal76, Van79]. Troesch [Sny79]. two [AGH79, BM76, BS76, GM77, Gra77, GMJ76, IKSN79, JS78, Lat75, MO78, Wam77, Wat78, WC79, Wil78, vdHSV79]. two-component [BS76]. two-dimensional [AGH79]. two-point [JS78, Wat78]. two-variable [MO78]. type [Ano75c, GC78, MO78, RT79, Sal76, Spi77]. types [Lat75].

Udine [dG77]. Uniform [Neu78].

University [Kau79]. unstable [Gra77]. updating [DS77]. USA [dG77]. Use [FK75]. useful [MV79]. using [Die75, DP77, Han79, MHH75, Rob77, Sal77a, Sas76, Sid76, Wuy75]. value [Aga79a, Gra77, JS78, SD76, WC79]. valued [Sta79]. values [Wat78]. variable [AM77, GMJ76, MO78]. variables [Mav79, Wam77]. Variational [Goo78c]. varying [Ano75a]. Veenman [Bul76]. Verlag [Goo78b, Goo78a, Kau79, dG77]. Vol [Kau79, dG77]. volume [Ano75b, Ano76b, Ano77b, Ano78].

W [Goo77b, Goo78b, Wat78]. Wageningen [Bul76]. weight [SD78]. Whittaker [SB78]. Wigert [Chi79]. window [Sid76]. workshop [Goo78b].

X [Ano76a, Goo77a]. XA [HMW77]. xiv [Kau79].

York [Kau79, dG77]. Young [Let76a].

zeroes [Nou77]. zeros [Nou75]. Zonen [Bul76].

References

Abbott:1978:EAD

[Abb78] James P. Abbott. An efficient algorithm for the determination of certain bifurcation points. Journal of Computational and Applied Mathematics, 4(1):19–27, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900153.

Abdrabbo:1979:NAP

[Abd79] N. A. Abdrabbo. A new approach to the problem of estimating spectral parameters of non-stationary time series models. Journal of Computational and Applied Mathematics, 5(2):125–129, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900081.

Aberth:1978:MEC

[Abe78] Oliver Aberth. A method for exact computation with rational numbers. Journal of Computational and Applied Mathematics, 4(4):285–288, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X7890027X.

Agarwal:1979:NSM

[Aga79a] Ravi P. Agarwal. The numerical solution of multipoint boundary value problems. Journal of Computational and Applied Mathematics, 5(1):17–24, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900226.

Agarwal:1979:PSN

[Aga79b] Ravi P. Agarwal. On periodic solutions of nonlinear second order differential systems. Journal of Computational and Applied Mathematics, 5(2):117–123, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X7990007X.

Avdelas:1979:TLE

[AGH79] G. Avdelas, S. Galanis, and A. Hadjidimos. Three-level extrapolated alternating direction implicit (E.A.D.I.) methods for the numerical solution of twodimensional second order elliptic problems. Journal of Computational and Applied Mathematics, 5(4):269–275, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900433.

Avdelas:1978:AEB

[AH78] G. Avdelas and A. Hadjidimos.

Analytical eigenvalue bounds for matrices arising when discretizing the self-adjoint second order elliptic partial differential equations. Journal of Computational and Applied Mathematics, 4(3):207–211, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900062.

Alt:1978:SOS

[Alt78] René Alt. A-stable one-step methods with step-size control for stiff systems of ordinary differential equations. Journal of Computational and Applied Mathematics, 4(1):29–35, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900165.

Andrade:1977:HAD

[AM77] Celia Andrade and S. McKee. High accuracy A.D.I. methods for fourth order parabolic equations with variable coefficients. Journal of Computational and Applied Mathematics, 3(1):11–14, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X77900195.

Anonymous:1975:BRA

[Ano75a] Anonymous. Book review: Adaptation of a linear system to a finite set of patterns occurring in an arbitrarily varying order: Teuvo Kohonen, Erkki Oja and Matti Ruohonen; Acta Polytechnica Scandinavica, Mathematics and Computing Machinery series No. 25. Journal of Computational and Applied Mathematics, 1(1):50, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900078.

Anonymous:1975:CV

[Ano75b] Anonymous. Contents of volume 1, 1975. Journal of Computational and Applied Mathematics, 1(4):289, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X75900212.

Anonymous:1975:SEP

[Ano75c] Anonymous. State estimation of polynomial type non-linear differential systems: Jussi orava; acta polytechnica scandinavica, mathematics and computing machinery series no. 25. Journal of Computational and Applied Mathematics, 1(4):246, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900145.

Anonymous:1976:BRT

[Ano76a] Anonymous. Book review: A theory for optimal PI-control of multivariable linear systems: Gustafsson, Sten E.: Acta Polytechnica Scandinavica, Mathematics and Computer Science Series No. 27. Helsinki 1975. 31 pp. ISBN 951-666-069-X. Journal of Computational and Applied Mathematics, 2(3):218, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900061.

Anonymous:1976:CV

[Ano76b] Anonymous. Contents of volume 2, 1976. Journal of Computational and Applied Mathematics, 2(4):282, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900516.

Anonymous:1976:EMC

[Ano76c] Anonymous. Erratum: A method for computing Feynman amplitudes with branch cuts [J. Comput. Appl. Math. 2(2) 73–76 (1976)]. Journal of Computational and Applied Mathematics, 2(3):218, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X76900073.

See [CGP76].

Anonymous:1977:CIS

[Ano77a] Anonymous. 1976 current index to statistics. Journal of Computational and Applied Mathematics, 3(4):298, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042777800265.

Anonymous:1977:CV

[Ano77b] Anonymous. Contents of volume 3, 1977. Journal of Computational and Applied Mathematics, 3(4):299, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042777800277.

Anonymous:1978:CV

[Ano78] Anonymous. Contents of volume 4, 1978. Journal of Computational and Applied Mathematics, 4(4):obc, December 1978.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X7890030X.

Ansari:1976:TAB

[Ans76] J. S. Ansari. A theorem on asymptotic behaviour of functional equations and its applications. Journal of Computational and Applied Mathematics, 2(1):35–39, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900371.

Ascher:1979:CMP

[AR79] U. Ascher and J. B. Rosen. A collocation method for parabolic quasilinear problems on general domains. Journal of Computational and Applied Mathematics, 5(3):183–191, September 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0377042779900049.

Axelsson:1978:SCA

[AS78] Owe Axelsson and Trond Steihaug. Some computational aspects in the numerical solution of parabolic equations. Journal of Computational and Applied Mathematics, 4(2):129–142, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900372.

Axelsson:1975:HOM

[Axe75] O. Axelsson. High-order methods for parabolic problems. Journal of Computational and Applied Mathematics, 1(1):5–16, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900029.

Breesch:1975:NNE

[BDG75] P. Breesch, J. De Kerf, and M. Goovaerts. A note on the numerical evaluation of integrals over strongly oscillating functions. Journal of Computational and Applied Mathematics, 1(1):47–49, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900066.

Benson:1979:NAS

[BE79] A. Benson and D. J. Evans. A normalised algorithm for the solution of symmetric general fiveterm linear systems of semibandwidth M. Journal of Computational and Applied Mathematics, 5(4):299–304, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X79900470.

Berkey:1978:PMA

[BF78] Dennis D. Berkey and Marvin I. Freedman. A perturbation method applied to the buckling of a compressed elastica. Journal of Computational and Applied Mathematics, 4(3):213–221, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900098.

Bekiroglu:1976:MAS

[BM76] Haluk Bekiroğlu and Keith L.

McRoberts. Mathematical analysis of a shuttle system based on two interdependent queues. Journal of Computational and Applied Mathematics, 2(3):195– 206, September 1976. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900048.

Bhattacharjee:1978:MII

[BM78] G. P. Bhattacharjee and K. L.

Majumder. Multivariate interval interpolation. Journal of Computational and Applied Mathematics, 4(4):295–300, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900293.

Brown:1977:REL

[BO77] D. R. Brown and M. J. O’Malley.

The role of eigenvalues in linear feature selection theory. Journal of Computational and Applied Mathematics, 3(4):243–249, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042777800150.

Branders:1975:ECC

[BP75] Maria Branders and Robert Piessens. An extension of Clenshaw–Curtis quadrature. Journal of Computational and Applied Mathematics, 1(1):55–65, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900091.

Brezinski:1976:CPA

[Bre76] Claude Brezinski. Computation of Padé approximants and continued fractions. Journal of Computational and Applied Mathematics, 2(2):113–123, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X76900152.

Bowman:1976:NAS

[BS76] K. O. Bowman and L. R. Shenton. A new algorithm for summing divergent series. Part 2: a two-component Borel summability model. Journal of Computational and Applied Mathematics, 2(4):259–266, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900486.

Bierbaum:1978:BIM

[BS78a] Fritz Bierbaum and Klaus-Peter Schwiertz. A bibliography on interval-mathematics. Journal of Computational and Applied Mathematics, 4(1):59–86, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900219.

Bowman:1978:ASS

[BS78b] K. O. Bowman and L. R. Shenton. Asymptotic series and Stieltjes continued fractions for a gamma function ratio. Journal of Computational and Applied Mathematics, 4(2):105–111, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900347.

Burniston:1978:SCA

[BS78c] E. E. Burniston and C. E. Siewert. On the solution of certain algebraic equations. Journal of Computational and Applied Mathematics, 4(1):37–39, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900177.

Bultheel:1976:BRP

[Bul76] A. Bultheel. Book review: Ph.

Th. Stol, Nonlinear parameter optimization: H. Veenman en Zonen B.V., Centre for Agricultural Publishing and Documentation, Wageningen 1975, 24 cm × 17 cm, 197 pages. Journal of Computational and Applied Mathematics, 2(1):70, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900437.

Bultheel:1979:RNL

[Bul79] A. Bultheel. Remark on “A new look at the Padé table and different methods for computing its elements” [2]. Journal of Computational and Applied Mathematics, 5(1):67, March 1979. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900305.

See [Cla75].

Barrodale:1978:FPS

[BW78] I. Barrodale and K. B. Wilson. A Fortran program for solving a nonlinear equation by Muller’s method. Journal of Computational and Applied Mathematics, 4(2):159–166, June 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900414.

Crane:1976:NSL

[CB76] C. M. Crane and D. M. Burley. Numerical studies of laminar flow in ducts and pipes. Journal of Computational and Applied Mathematics, 2(2):95–111, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900140.

Chatterjee:1976:NMM

[CD76] Anadi Kumar Chatterjee and Lokenath Debnath. Nonlinear mathematical model of the junction of tidal rivers. Journal of Computational and Applied Mathematics, 2(1):13–18, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900334.

Chatterjee:1978:NMM

[CD78] Anadi Kumar Chatterjee and Lokenath Debnath. Nonlinear mathematical model of inhomogeneous tidal rivers. Journal of Computational and Applied Mathematics, 4(3):223–227, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900104.

Cook:1977:FSO

[CG77] A. E. Cook and R. D. Gibson.

Fifth and sixth order PECE algorithms based on g-splines with improved stability characteristics. Journal of Computational and Applied Mathematics, 3(2):85–87, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X7790002X.

Chisholm:1976:MCF

[CGP76] J. S. R. Chisholm, A. C. Genz, and M. Pusterla. A method for computing Feynman amplitudes with branch cuts. Journal of Computational and Applied Mathematics, 2(2):73–76, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900103.

See erratum [Ano76c].

Chihara:1979:GSW

[Chi79] T. S. Chihara. On generalized Stieltjes–Wigert and related orthogonal polynomials. Journal of Computational and Applied Mathematics, 5(4):291–297, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X79900469.

Chan:1975:PCS

[CK75] Chorkin Chan and Melvin Klassen. A performance comparison study between subroutine packages LIN-SYS, IBMSSP and IMSL for solving systems of linear equations. Journal of Computational and Applied Mathematics, 1(2):111–113, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900285.

Corliss:1977:CST

[CL77] George Corliss and David Lowery. Choosing a stepsize for Taylor series methods for solving ODE’S. Journal of Computational and Applied Mathematics, 3(4):251–256, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800162.

Claessens:1975:NLP

[Cla75] G. Claessens. A new look at the Padé table and the different methods for computing its elements. Journal of Computational and Applied Mathematics, 1(3):141–152, September 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900327.

See remark [Bul79].

Coleman:1978:RBA

[Col78]  John P. Coleman.  A “re-

fined Born approximation” and Aitken’s Δ^{2} method. Journal of Computational and Applied Mathematics, 4(2):101–103, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900335.

Claessens:1979:CNN

[CW79] G. Claessens and L. Wuytack. On the computation of non-normal Padé approximants. Journal of Computational and Applied Mathematics, 5(4):283–289, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900457.

Dahiya:1977:CDL

[Dah77] R. S. Dahiya. Computation of n-dimensional Laplace transforms. Journal of Computational and Applied Mathematics, 3(3): 185–188, September 1977. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704277780006X.

Danloy:1976:CSH

[Dan76] B. Danloy. On the choice of signs for Householder’s matrices. Journal of Computational and Applied Mathematics, 2(1):67–69, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900425.

deBruin:1978:ACG

[dB78] M. G. de Bruin. An algorithm for calculating generalized continued fractions. Journal of Computational and Applied Mathematics, 4(3):177–179, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900013.

DePril:1976:BCT

[DDG76] N. De Pril, L. D’Hooge, and M. J. Goovaerts. A bibliography on credibility theory and its applications. Journal of Computational and Applied Mathematics, 2(1):55–62, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900401.

DeKerf:1975:BFS

[De 75a] Joseph L. F. De Kerf. A bibliography on fuzzy sets. Journal of Computational and Applied Mathematics, 1(3):205–212, September 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900364.

DeKerf:1975:IMS

[De 75b] Joseph L. F. De Kerf. The interpolation method of Sprague–Karup. Journal of Computational and Applied Mathematics, 1(2):101–110, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900273.

DeMeersman:1975:MLS

[De 75c] R. De Meersman. A method for least squares solution of systems with a cyclic rectangular coefficient matrix. Journal of Computational and Applied Mathematics, 1(1):51–54, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X7590008X.

DeMeersman:1976:DLA

[De 76] R. De Meersman. A derivation of the Levinson algorithm for solving linear systems with symmetric positive definite Toeplitz matrix. Journal of Computational and Applied Mathematics, 2(1):63–65, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900413.

DeVylder:1977:NPK

[De 77] Fl. De Vylder. A new proof for a known result in risk theory. Journal of Computational and Applied Mathematics, 3(4):277–279, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800216.

DeMeersman:1979:LSN

[De 79] R. De Meersman. Least squares with non-linear equality constraints application to closing of balances. Journal of Computational and Applied Mathematics, 5(4):277–281, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900445.

Dekker:1977:GRK

[Dek77] K. Dekker. Generalized Runge– Kutta methods for coupled systems of hyperbolic differential equations. Journal of Computational and Applied Mathematics, 3(4):221–233, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800137.

Devooght:1976:ACR

[Dev76] J. Devooght. Analytic continuation by reproducing kernel methods combined with Padé approximations. Journal of Computational and Applied Mathematics, 2(4):231–239, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900449.

Dey:1977:PIS

[Dey77] S. K. Dey. Perturbed iterative solution of nonlinear equations with applications to fluid dynamics. Journal of Computational and Applied Mathematics, 3(1):15–30, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900201.

Dey:1978:PIS

[Dey78] S. K. Dey. Perturbed iterative solution of coupled nonlinear systems with application to fluid dynamics. Journal of Computational and Applied Mathematics, 4(2):117–128, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900360.

deGroot:1977:BRM

[dG77] R. de Groot. Book review: Mathematical systems theory: Lecture notes in economics and mathematical systems. Vol. 131. Proceedings of the International Symposium, Udine, Italy, June 1975, Marchesini G., Padua, Italy and Mitter, S. K., Massachusetts Institute of Technology, USA, editors. Springer Verlag, New York 1976, 408 pages. Journal of Computational and Applied Mathematics, 3(1):75–76, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900249.

Dierckx:1975:ASD

[Die75] P. Dierckx. An algorithm for smoothing, differentiation and integration of experimental data using spline functions. Journal of Computational and Applied Mathematics, 1(3):165–184, September 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900340.

Dierckx:1977:ALS

[Die77] P. Dierckx. An algorithm for leastsquares fitting of cubic spline surfaces to functions on a rectilinear mesh over a rectangle. Journal of Computational and Applied Mathematics, 3(2):113–129, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900079.

DeDoncker:1976:BAI

[DP76] E. De Doncker and R. Piessens.

A bibliography on automatic integration. Journal of Computational and Applied Mathematics, 2(4):273–279, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900504.

Dierckx:1977:CFC

[DP77] P. Dierckx and R. Piessens. Calculation of Fourier coefficients of discrete functions using cubic splines. Journal of Computational and Applied Mathematics, 3(3):207–209, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800101.

DeMeersman:1977:UFL

[DS77] R. De Meersman and P. Siau.

On updating formulas for least squares solutions and pseudoinverses. Journal of Computational and Applied Mathematics, 3(4):293–297, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S037704277780023X.

Dunham:1979:CAE

[Dun79a] Charles B. Dunham. Chebyshev approximation by exponentialpolynomial sums. Journal of Computational and Applied Mathematics, 5(1):53–57, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900275.

Dunham:1979:SPP

[Dun79b] Charles B. Dunham. A survey of published programs for best approximation. Journal of Computational and Applied Mathematics, 5(2):141–143, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X7990010X.

Ebert:1977:COR

[Ebe77] Udo Ebert. Computation of optimal radiation treatment plans. Journal of Computational and Applied Mathematics, 3(2):99–104, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900043.

Evans:1977:LMN

[EF77] D. J. Evans and S. O. Fatunla.

A linear multistep numerical integration scheme for solving systems of ordinary differential equations with oscillatory solutions. Journal of Computational and Applied Mathematics, 3(4):235–241, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800149.

Evans:1979:IHT

[EG79] N. T. S. Evans and A. R. Gourlay.

Implementation of the “Hopscotch” technique in the numerical solution of diffusion problems. Journal of Computational and Applied Mathematics, 5(1):47–52, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900263.

Einarsson:1977:SSL

[Ein77] Bo Einarsson. Solving a system of linear inequalities for generating test data. Journal of Computational and Applied Mathematics, 3(4):269–271, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800198.

Einarsson:1979:BEN

[Ein79] Bo Einarsson. Bibliography on the evaluation of numerical software.

Journal of Computational and Applied Mathematics, 5(2):145–159, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900111.

Evans:1977:CEE

[Eva77] David J. Evans. Computation of eigenvalues and eigenvectors of a symmetric quindiagonal matrix. Journal of Computational and Applied Mathematics, 3(2):131–141, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900080.

Freedman:1975:UNL

[FK75] Marvin I. Freedman and James L. Kaplan. Use of a non-linear clock in the perturbation analysis of time optimal control problems. Journal of Computational and Applied Mathematics, 1(4): 219–228, December 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X75900108.

Greenspan:1978:CSP

[GC78] Donald Greenspan and John Collier. Computer studies of planetary-type evolution. Journal of Computational and Applied Mathematics, 4(4):235–257, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900220.

Genz:1977:NMA

[Gen77] Alan C. Genz. A nonlinear method for the acceleration of the convergence of multidimensional sequences. Journal of Computational and Applied Mathematics, 3(3):181–184, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800058.

Gretton:1979:CTC

[GJ79] H. W. Gretton and K. A. Jukes.

Computations of Tauberian constants connected with equiconvergence and the quadratfrei numbers. Journal of Computational and Applied Mathematics, 5(1):25–27, March 1979. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900238.

Gourlay:1977:CHM

[GM77] A. R. Gourlay and S. McKee. The construction of hopscotch methods for parabolic and elliptic equations in two space dimensions with a mixed derivative. Journal of Computational and Applied Mathematics, 3(3):201–206, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800095.

Gourlay:1978:LOD

[GM78] A. R. Gourlay and J. Ll. Morris. Locally one-dimensional hopscotch methods. Journal of Computational and Applied Mathematics, 4(4):269–273, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900256.

Gourlay:1979:HME

[GM79] A. R. Gourlay and S. McKee. Hopscotch methods for elliptic partial differential equations. Journal of Computational and Applied Mathematics, 5(2):103–110, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X79900056.

Graves-Morris:1976:ATV

[GMJ76] P. R. Graves-Morris and R. Hughes Jones. An analysis of two variable rational approximants. Journal of Computational and Applied Mathematics, 2(1):41–48, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900383.

Goovaerts:1977:BRO

[Goo77a] M. J. Goovaerts. Book review: Optimal controller design by nonlinear and game theoretic method, Hämäläinen, Raimo P., in: Acta Polytechnica Scandinavica, Mathematics and Computer Science Series, No. 28 (1976), ISSN 0355-2713, ISBN 951-666-086-X. Journal of Computational and Applied Mathematics, 3(4):298, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800253.

Goovaerts:1977:BRP

[Goo77b] M. J. Goovaerts. Book review:

Preservation of infinite divisibility under mixing and related topics, 33, F. W. Steutel, in: Mathematical Centre Tracts. Stichting Mathematisch Centrum (1977), 1–99. Journal of Computational and Applied Mathematics, 3(4):298, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800241.

Goovaerts:1978:BRI

[Goo78a] M. J. Goovaerts. Book review: An introduction to the regerative method for simulation analysis: M. A. Crane, A. J. Lemoine, Berlin, Springer-Verlag, 1977, 24 cm × 17 cm. 111 p., price DM 18. Journal of Computational and Applied Mathematics, 4(3):212, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900074.

Goovaerts:1978:BRD

[Goo78b] M. J. Goovaerts. Book review: Differential games and applications: Proceedings of a workshop, Enschede, 1977: P. Hagedorn, H. W. Knobloch, G. J. Olsder, editors. Berlin, Springer-Verlag, 1977, 17 cm × 24 cm, 236 p., 60 fig., 6 tab., price DM 24,80. Journal of Computational and Applied Mathematics, 4(4):258, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900232.

Goovaerts:1978:BRV

[Goo78c] M. J. Goovaerts. Book review: Variational methods in mathematics, science and engineering: Karel Rektorys, (1977), 516 pp., D. Reidel Publishing Company. Journal of Computational and Applied Mathematics, 4(3):198, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X78900049.

Graney:1977:NSU

[Gra77] L. Graney. The numerical solution of unstable two point boundary value problems. Journal of Computational and Applied Mathematics, 3(3):173–180, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042777800046.

Greenspan:1977:CSI

[Gre77] Donald Greenspan. Computer studies of interactions of particles with differing masses. Journal of Computational and Applied Mathematics, 3(3):145–154, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800010.

Goovaerts:1978:BET

[GV78] M. J. Goovaerts and P. Van Goethem. On a Berry–Esseen theorem for compound Poisson processes. Journal of Computational and Applied Mathematics, 4(2):93–100, June 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900323.

Hanson:1979:CIP

[Han79] Richard J. Hanson. Computing an interior point for inequalities using linear optimization. Journal of Computational and Applied Mathematics, 5(1):69–70, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900317.

Hennessy:1978:CAL

[Hen78] T. R. Hennessy. A comparative approach to Laplace transform inversion. Journal of Computational and Applied Mathematics, 4(2):89–91, June 1978.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900311.

Hautot:1979:CES

[HM79] André Hautot and Alphonse Magnus. Calculation of the eigenvalues of Schrödinger equations by an extension of Hill’s method. Journal of Computational and Applied Mathematics, 5(1):3–15, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900214.

Hoskins:1977:RCM

[HMW77] W. D. Hoskins, D. S. Meek, and D. J. Walton. A rapidly convergent method for the solution of of the matrix equation XA + AY=F. Journal of Computational and Applied Mathematics, 3(3):211–215, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800113.

Houstis:1978:CNM

[Hou78] E. N. Houstis. The complexity of numerical methods for elliptic partial differential equations. Journal of Computational and Applied Mathematics, 4(3): 191–197, September 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900037.

Ikeuchi:1979:SRJ

[IKSN79] Masatoshi Ikeuchi, Hiroshi Kobayashi, Hideo Sawami, and Hiroshi Niki. On the spectral radius of the Jacobi iteration matrix for a rectangular region with two different media. Journal of Computational and Applied Mathematics, 5(4):247–258, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900408.

Iserles:1978:AEF

[Ise78] Arieh Iserles. A-acceptable exponentially fitted combinations of three Padé approximations. Journal of Computational and Applied Mathematics, 4(2):143–146, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900384.

Jeffreson:1978:LSC

[JC78] Carl P. Jeffreson and Ee-Pan Chow. Least squares coefficients for a quadrature formula for Laplace transform inversion. Journal of Computational and Applied Mathematics, 4(1):53–58, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900207.

Johnson:1977:ILS

[Joh77] R. C. Johnson. Interpolation by local space curves. Journal of Computational and Applied Mathematics, 3(2):79–84, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900018.

Jain:1978:HOD

[JS78] M. K. Jain and J. S. V. Saldanha. Higher order difference methods for second order two-point boundary-value problems. Journal of Computational and Applied Mathematics, 4(3): 199–206, September 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900050.

Kaufman:1979:BRI

[Kau79] L. Kaufman. Book review: Integer programming and related areas — A classified bibliography 1976– 1978: Compiled at the Institut für Ökonometrie und Operations Research, University of Bonn. Edited by Dirk Hausman. 1978, xiv + 314 pages. Soft cover, DM 32, US 16. Springer-Verlag, Berlin, Heidelberg, New York. Vol. 160 of “Lecture Notes in Economics and Mathematical Systems”, Editors M. Beckmann and H. P. Künzi. Journal of Computational and Applied Mathematics, 5(2):102, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900044.

Kloek:1975:PEI

[KL75] T. Kloek and F. B. Lempers. Postexperiment introduction of constraints on parameters. Journal of Computational and Applied Mathematics, 1(4):235–237, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900121.

Klein:1979:SFS

[Kle79] G. Klein. On spline functions and statistical regularization of ill-posed problems. Journal of Computational and Applied Mathematics, 5(4):259–264, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X7990041X.

Kammler:1978:BAE

[KM78] David W. Kammler and Robert J.

McGlinn. A bibliography for approximation with exponential sums. Journal of Computational and Applied Mathematics, 4(2):167–173, June 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900426.

Laporte:1978:SPT

[Lap78] Gilbert Laporte. The seriation problem and the travelling salesman problem. Journal of Computational and Applied Mathematics, 4(4):259–268, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900244.

Latouche:1975:OAP

[Lat75] Guy Latouche. Optimal allocation of priority in a M/M/1 queue with two types of customers. Journal of Computational and Applied Mathematics, 1(2):85–91, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X7590025X.

Lether:1976:BYQ

[Let76a] Frank Lether. On Birkhoff–Young quadrature of analytical functions. Journal of Computational and Applied Mathematics, 2(2):81–84, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900127.

Lether:1976:CDI

[Let76b] Frank G. Lether. Computation of double integrals over a triangle. Journal of Computational and Applied Mathematics, 2(3):219–224, September 1976. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900085.

Lether:1977:MQF

[Let77]  Frank G. Lether.  Modified

quadrature formulas for functions with nearby poles. Journal of Computational and Applied Mathematics, 3(1):3–9, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X77900171.

Lether:1978:CGL

[Let78] Frank G. Lether. On the construction of Gauss–Legendre quadrature rules. Journal of Computational and Applied Mathematics, 4(1):47–52, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900190.

Levin:1975:NIL

[Lev75] David Levin. Numerical inversion of the Laplace transform by accelerating the convergence of Bromwick’s integral. Journal of Computational and Applied Mathematics, 1(4):247–250, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900157.

Lewanowicz:1979:CRR

[Lew79] Stanislaw Lewanowicz. Construction of a recurrence relation for modified moments. Journal of Computational and Applied Mathematics, 5(3):193–206, September 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0377042779900050.

Litinsky:1978:BAN

[Lit78] E. G. Litinsky. Best approximation and numerical methods in solution of differential equations. Journal of Computational and Applied Mathematics, 4(2):113–116, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900359.

Longman:1975:ABR

[Lon75] I. M. Longman. Application of best rational function approximation for Laplace transform inversion. Journal of Computational and Applied Mathematics, 1(1):17–23, March 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X75900030.

Laurie:1979:CGQ

[LR79] Dirk P. Laurie and Laurette Rolfes. Computation of Gaussian quadrature rules from modified moments. Journal of Computational and Applied Mathematics, 5(3):235–243, September 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0377042779900086.

Luke:1976:CER

[Luk76] Yudell L. Luke. Chebyshev expansions and rational approximations. Journal of Computational and Applied Mathematics, 2(2):85–93, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900139.

Mavrides:1979:NPC

[Mav79] Lazaros P. Mavrides. Nonlinear programming with cumulatively bounded variables. Journal of Computational and Applied Mathematics, 5(3):163–169, September 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0377042779900013.

Meeusen:1975:CDD

[MC75] W. Meeusen and L. Cuyvers.

Clique detection in directed graphs: a new algorithm. Journal of Computational and Applied Mathematics, 1(3):185– 203, September 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X75900352.

Meersman:1975:STA

[Mee75] Robert Meersman. A survey of techniques in applied computational complexity. Journal of Computational and Applied Mathematics, 1(1):39–46, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900054.

Melka:1978:MCA

[Mel78] Richard F. Melka. The method of coordinate averaging. Journal of Computational and Applied Mathematics, 4(1):41–45, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900189.

Melka:1979:IP

[Mel79] Richard F. Melka. Implicit perturbations. Journal of Computational and Applied Mathematics, 5(3):177–182, September 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0377042779900037.

Mund:1975:AIF

[MHH75] E. H. Mund, P. Hallet, and J. P. Hennart. An algorithm for the interpolation of functions using quintic splines. Journal of Computational and Applied Mathematics, 1(4):279–288, December 1975. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900200.

Miklosko:1975:RCD

[Mik75] Jozef Miklosko. A recursive computation of the determinant of a pentadiagonal matrix. Journal of Computational and Applied Mathematics, 1(2):73–78, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900236.

Miklosko:1977:ACC

[Mik77] Jozef Miklosko. An algorithm for calculating continued fractions. Journal of Computational and Applied Mathematics, 3(4): 273–275, December 1977. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0377042777800204.

Mizrahi:1975:GHPa

[Miz75a] Maurice M. Mizrahi. Generalized Hermite polynomials. Journal of Computational and Applied Mathematics, 1(3):137–140, September 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900315.

Mizrahi:1975:GHPb

[Miz75b] Maurice M. Mizrahi. Generalized Hermite polynomials. Journal of Computational and Applied Mathematics, 1(4):273–277, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900194.

Murphy:1978:TVG

[MO78] J. A. Murphy and M. R. O’Donohoe. A two-variable generalization of the Stieltjes-type continued fraction. Journal of Computational and Applied Mathematics, 4(3):181–190, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900025.

Meire:1979:URC

[MV79] R. Meire and A. Vanderbauwhede. A useful result for certain linear periodic ordinary differential equations. Journal of Computational and Applied Mathematics, 5(1):59–61, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900287.

Neuts:1977:ABC

[Neu77] M. F. Neuts. An annoted bibliography on computational probability: Part 1. Journal of Computational and Applied Mathematics, 3(1):53–74, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900237.

Neuman:1978:UAS

[Neu78] E. Neuman. Uniform approximation by some Hermite interpolating splines. Journal of Computational and Applied Mathematics, 4(1):7–9, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X7890013X.

Nerinckx:1976:CNL

[NH76] D. Nerinckx and A. Haegemans. A comparison of nonlinear equation solvers. Journal of Computational and Applied Mathematics, 2(2):145–148, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900176.

Nourein:1975:IFS

[Nou75] Abdel-Wahab M. Nourein. An iteration formula for the simultaneous determination of the zeros of a polynomial. Journal of Computational and Applied Mathematics, 1(4):251–254, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900169.

Nourein:1977:INM

[Nou77] A. W. Nourein. An improvement on Nourein’s method for the simultaneous determination of the zeroes of a polynomial (an algorithm). Journal of Computational and Applied Mathematics, 3(2):109–112, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X77900067.

Niethammer:1975:RSM

[NS75] W. Niethammer and J. Schade. On a relaxed SOR-method applied to nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 1(3): 133–136, September 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X75900303.

Oliver:1977:SRE

[Oli77] J. Oliver. On the sensitivity to rounding errors of Chebyshev series approximations. Journal of Computational and Applied Mathematics, 3(2):89–98, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900031.

Oliver:1979:REP

[Oli79] J. Oliver. Rounding error propagation in polynomial evaluation schemes. Journal of Computational and Applied Mathematics, 5(2):85–97, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900020.

OShea:1978:ASS

[O’S78]  Brendan B. O’Shea.  Algo-

rithms for the solution of systems of coupled second-order ordinary differential equations. Journal of Computational and Applied Mathematics, 4(1):11–17, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900141.

Paulik:1976:OAB

[Pau76] Augustin Paulik. On the optimal approximation of bounded linear functionals in Hilbert spaces with inner product invariant in rotation or translation. Journal of Computational and Applied Mathematics, 2(4):267–272, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900498.

Piessens:1975:COI

[PB75] Robert Piessens and Maria Branders. Computation of oscillating integrals. Journal of Computational and Applied Mathematics, 1(3):153–164, September 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900339.

Piessens:1976:BNI

[PD76] R. Piessens and N. D. P. Dang. A bibliography on numerical inversion of the Laplace transform and applications: a supplement. Journal of Computational and Applied Mathematics, 2(3):225–228, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900097.

Peschanel:1975:ACQ

[Pes75] F. D. Peschanel. An approach to a consistent qualification of decision procedures in probabilistic information processing. Journal of Computational and Applied Mathematics, 1(4):239–245, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900133.

Piessens:1975:CFD

[PH75] Robert Piessens and Ann Haegemans. Cubature formulas of degree eleven for symmetric planar regions. Journal of Computational and Applied Mathematics, 1(2):79–83, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X75900248.

Picca:1976:LTA

[Pic76] D. Picca. Laplace-transform approach to the quantal Coulomb problem. Journal of Computational and Applied Mathematics, 2(1):49–54, ???? 1976. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900395.

Piessens:1975:BNI

[Pie75] Robert Piessens. A bibliography on numerical inversion of the Laplace transform and applications. Journal of Computational and Applied Mathematics, 1(2):115–128, June 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X75900297.

Pindor:1976:SAC

[Pin76] M. Pindor. A simplified algorithm for calculating the Padé table derived from Baker and Longman schemes. Journal of Computational and Applied Mathematics, 2(4):255–257, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900474.

Poleunis:1977:CSA

[PL77] F. Poleunis and O. Leroy. The construction of self-adjoint operators of Hill with periodic eigenfunctions. Journal of Computational and Applied Mathematics, 3(3):189–194, September 1977. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042777800071.

Paiano:1975:ITC

[PP75] G. Paiano and D. Picca. Integral transforms as computational tools in quantum mechanics. Journal of Computational and Applied Mathematics, 1(2):93–100, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X75900261.

Piessens:1975:E

[PW75] Robert Piessens and Luc Wuytack. Editorial. Journal of Computational and Applied Mathematics, 1(1):3, March 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X75900017.

Papamichael:1977:LE

[PW77] N. Papamichael and J. R. Whiteman. Letter to the editors. Journal of Computational and Applied Mathematics, 3(1):10, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X77900183.

Rahman:1978:NSF

[Rah78] Matiur Rahman. On the numerical solution of the flow between a rotating and a stationary disk. Journal of Computational and Applied Mathematics, 4(4):289–293, December 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900281.

Rayski:1977:RBA

[Ray77] Jerzy Rayski. A refined Born approximation. Journal of Computational and Applied Mathematics, 3(1):31–34, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X77900213.

Razzaghi:1976:JCO

[Raz76] M. Razzaghi. On the Jacobi condition for optimizing an engineering problem. Journal of Computational and Applied Mathematics, 2(2):77–80, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900115.

Razzaghi:1978:PAS

[Raz78] M. Razzaghi. On the pole assignment for single-input systems with time-delays. Journal of Computational and Applied Mathematics, 4(2):155–157, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900402.

Rabbo:1978:CTD

[RM78] N. A. Abd Rabbo and A. Mattar. A CUSUM test for detecting change in the transfer functions of open loop stochastic systems. Journal of Computational and Applied Mathematics, 4(2):147–154, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900396.

Roberts:1977:ADP

[Rob77] D. E. Roberts. An analysis of double power series using rotationally covariant approximants. Journal of Computational and Applied Mathematics, 3(4):257–262, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042777800174.

Robinson:1979:CNI

[Rob79] Ian Robinson. A comparison of numerical integration programs. Journal of Computational and Applied Mathematics, 5(3):207– 223, September 1979. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0377042779900062.

Rodabaugh:1979:LOS

[RT79] D. J. Rodabaugh and S. Thompson. Low-order A_{0}-stable Adamstype correctors. Journal of Computational and Applied Mathematics, 5(3):225–233, September 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0377042779900074.

Roseman:1976:CNS

[RZ76] Joseph Roseman and Gideon Zwas. On the correct numerical shock speed. Journal of Computational and Applied Mathematics, 2(1):3–6, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X76900310.

Reddy:1975:NED

[RZH75] P. J. Reddy, H.-J. Zimmermann, and Asghar Husain. Numerical experiments on DFP-method, a powerful function minimization technique. Journal of Computational and Applied Mathematics, 1(4):255–265, December 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900170.

Sack:1979:NMS

[Sac79] R. A. Sack. A numerical method for simultaneous convergence to the elements of a sequence. Journal of Computational and Applied Mathematics, 5(1):29–35, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X7990024X.

Salzer:1976:NFT

[Sal76] Herbert E. Salzer. A new form of trigonometric orthogonality and Gaussian-type quadrature. Journal of Computational and Applied Mathematics, 2(4):241–248, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900450.

Salzer:1977:ILT

[Sal77a] Herbert E. Salzer. Inverse Laplace transforms of osculatory and hyperosculatory interpolation polynomials using “near-Chebyshev” nodes. Journal of Computational and Applied Mathematics, 3(3):217, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042777800125.

Salzer:1977:MED

[Sal77b] Herbert E. Salzer. Minimal error difference formulas. Journal of Computational and Applied Mathematics, 3(4):263–267, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042777800186.

Sastry:1976:FDA

[Sas76] S. S. Sastry. Finite difference approximations to one-dimensional parabolic equations using a cubic spline technique. Journal of Computational and Applied Mathematics, 2(1):23–26, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900358.

Shenton:1976:NAS

[SB76a] L. R. Shenton and K. O. Bowman.

A new algorithm for summing divergent series. Part 1. Basic theory and illustrations. Journal of Computational and Applied Mathematics, 2(3):151–167, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900012.

Siewert:1976:EAS

[SB76b] C. E. Siewert and E. E. Burniston. An exact analytical solution of x coth x = αx^{2} + 1. Journal of Computational and Applied Mathematics, 2(1):19–21, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900346.

Shenton:1977:NAS

[SB77] L. R. Shenton and K. O. Bowman. A new algorithm for summing divergent series: Part 3: Applications. Journal of Computational and Applied Mathematics, 3(1):35–51, March 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X77900225.

Shenton:1978:GCF

[SB78] L. R. Shenton and K. O. Bowman. Generalized continued fractions and Whittaker’s approach. Journal of Computational and Applied Mathematics, 4(1):3–5, ???? 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X78900128.

Sanchez-Dehesa:1976:CHM

[SD76] Jesus Sánchez-Dehesa. On the conditions for a Hamiltonian matrix to have an eigen-value density with some prescribed characteristics. Journal of Computational and Applied Mathematics, 2(4):249–254, December 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0771050X76900462.

Sanchez-Dehesa:1977:ASJ

[SD77]  Jesus Sánchez-Dehesa.  The

asymptotical spectrum of Jacobi matrices. Journal of Computational and Applied Mathematics, 3(3):167–171, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042777800034.

Sanchez-Dehesa:1978:SJM

[SD78] Jesus Sanchez-Dehesa. The spectrum of Jacobi matrices in terms of its associated weight function. Journal of Computational and Applied Mathematics, 4(4): 275–283, December 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X78900268.

Sanchez-Dehesa:1979:GSO

[SD79] Jesus Sánchez-Dehesa. On a general system of orthogonal qpolynomials. Journal of Computational and Applied Mathematics, 5(1):37–45, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900251.

Sidi:1975:CCP

[Sid75a] Avram Sidi. Computation of the Chebyshev–Padé table. Journal of Computational and Applied Mathematics, 1(2):69–71, June 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900224.

Sidi:1975:ASI

[Sid75b] Avram Sidi. On the approximation of square-integrable functions by exponential series. Journal of Computational and Applied Mathematics, 1(4):229–234, December 1975. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X7590011X.

Sidi:1976:BRF

[Sid76] Avram Sidi. Best rational function approximation to Laplace transform inversion using a window function. Journal of Computational and Applied Mathematics, 2(3):187–194, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900036.

Snyman:1979:CDN

[Sny79]  J. A. Snyman.  Continuous

and discontinuous numerical solutions to the Troesch problem. Journal of Computational and Applied Mathematics, 5(3): 171–175, September 1979. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0377042779900025.

Siewert:1979:S

[SP79] C. E. Siewert and J. S. Phelps III. On the solution of a tan(ζ − kπ) + tanh ζ = 0. Journal of Computational and Applied Mathematics, 5(2):99–101, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900032.

Spitzer:1977:NSE

[Spi77] John J. Spitzer. A numerically stable and efficient technique for the maintenance of positive definiteness in the Hessian for Newton-type methods. Journal of Computational and Applied Mathematics, 3(2):105–108, June 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X77900055.

Strom:1976:DOT

[SS76] Torsten Ström and Per Svensson. On the determination of an orthogonal transformation. Journal of Computational and Applied Mathematics, 2(1):7–11, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900322.

Starkand:1979:EFM

[Sta79] Yair Starkand. Explicit formulas for matrix-valued Padé approximants. Journal of Computational and Applied Mathematics, 5(1):63–66, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900299.

Stetter:1975:PPS

[SU75] Hans J. Stetter and Christoph W.

Ueberhuber. Proposal for population studies in numerical quadrature. Journal of Computational and Applied Mathematics, 1(3): 213–215, September 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X75900376.

Sloan:1977:RTM

[SW77] D. M. Sloan and G. Wilks. The Riccati transformation method and the computation of eigenvalues of complex linear differential systems. Journal of Computational and Applied Mathematics, 3(3):195–199, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800083.

Teugels:1976:BSM

[Teu76] Jozef L. Teugels. A bibliography on semi-Markov processes. Journal of Computational and Applied Mathematics, 2(2):125–144, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900164.

VandeVel:1979:HIT

[Van79] H. Van de Vel. Haar intervals for trigonometric polynomials. Journal of Computational and Applied Mathematics, 5(4):265–267, December 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900421.

vanDooren:1976:AAN

[vDdR76] Paul van Dooren and Luc de Ridder. An adaptive algorithm for numerical integration over an ndimensional cube. Journal of Computational and Applied Mathematics, 2(3):207–217, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X7690005X.

vanderHouwen:1979:CTI

[vdHSV79] P. J. van der Houwen, B. P. Sommeijer, and J. G. Verwer. Comparing time integrators for parabolic equations in two space dimensions with a mixed derivative. Journal of Computational and Applied Mathematics, 5(2):73–83, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900019.

Verwer:1977:CST

[Ver77] J. G. Verwer. A class of stabilized three-step Runge–Kutta methods for the numerical integration of parabolic equations. Journal of Computational and Applied Mathematics, 3(3):155–166, September 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800022.

Wu:1975:NMA

[WA75] C. L. Wu and R. J. Adler. Nonlinear matrix algebra and engineering applications. Part 1: Theory and linear form matrix. Journal of Computational and Applied Mathematics, 1(1):25–37, March 1975. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X75900042.

Wu:1976:NMA

[WA76] C. L. Wu and R. J. Adler. Nonlinear matrix algebra and engineering applications. Part 2: Polynomial form matrices. Journal of Computational and Applied Mathematics, 2(3):169–185, September 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X76900024.

Wambecq:1976:NMS

[Wam76] A. Wambecq. Nonlinear methods in solving ordinary differential equations. Journal of Computational and Applied Mathematics, 2(1):27–33, ???? 1976. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X7690036X.

Wambecq:1977:PFT

[Wam77] A. Wambecq. Plotting a function of two variables. Journal of Computational and Applied Mathematics, 3(4):281–291, December 1977. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0377042777800228.

Waterschoot:1978:BRN

[Wat78] W. Waterschoot. Book review:

A numerical study of stiff twopoint boundary values, P. W. Hemker: Mathematical Centre Tracts 80. Edited by Stichting Mathematisch Centrum, Amsterdam, 1977; 24 cm × 16 cm; 178 blz; price f 22. Journal of Computational and Applied Mathematics, 4(3):212, September 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X78900086.

Watson:1979:FPM

[Wat79] Layne T. Watson. Fixed points of C^{2} maps. Journal of Computational and Applied Mathematics, 5(2):131–139, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0771050X79900093.

Westreich:1979:NMD

[WC79] David Westreich and Baruch Cahlon. Numerical methods for discontinuous linear two point boundary value problems. Journal of Computational and Applied Mathematics, 5(2):111–115, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0771050X79900068.

Wilhelmsen:1978:DTG

[Wil78] Don R. Wilhelmsen. Degree two Gaussian cubatures in the plane. Journal of Computational and Applied Mathematics, 4(3): 229–232, September 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X78900116.

Wuytack:1975:NIU

[Wuy75] L. Wuytack. Numerical integration by using nonlinear techniques. Journal of Computational and Applied Mathematics, 1(4): 267–272, December 1975. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X75900182.

Wuytack:1979:EN

[Wuy79] L. Wuytack. Editorial note. Journal of Computational and Applied Mathematics, 5(1):2, March 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0771050X79900202.